5.) A car traveling at 15 m/s passes a van that is starting from rest with an acceleration of 3 m/s². (a) How long will it take the van to catch up with the car? (b) How far has the car traveled when the van catches up with it?
1
Expert's answer
2019-08-26T10:41:49-0400

As the car moves with the constant speed, the movement can be described by the following equation:

"S = V*t"

As the van moves with the constant acceleration, the movement can be described by the following equation:

"S = V_0 *t + at^2\/2" , where V0 = 0 (m/s)

If the van catches up with the car then they have travelled the same distance for the same time:

S1 = S2; t1 = t2

Find time needed to catch up with the car:

"V*t = at^2\/2"

"2Vt = at^2"

"at^2-2Vt = 0"

"t(at-2V)=0"

t = 0 -- the moment when the car passes the van;

"at-2V=0"

"at = 2V"

"t = 2V\/a"

"t = 2*15\/3 = 10 (s)"

Find the distance, the car has travelled when the van catches up with it:

"S = V*t"

"S = 15 * 10 = 150 (m)"

Answer: it takes 10 seconds for the van to catch up with the car; at the moment of catching up the car has travelled 150 m.